OpenStudy (anonymous):

How many square inches are in 3 square feet?

OpenStudy (anonymous):

the answer is 432 but i don't know how they got that answer

OpenStudy (anonymous):

1 foot =12 inches

OpenStudy (anonymous):

a square foot means, we square the foot, so (foot)^2=(12in)^2=144 in^2

OpenStudy (anonymous):

so in 3 (foot)^2=3*144in^2=432in^2